OpenStudy (anonymous):

Solve the system using the addition method. 3x + 4y = -10 5x - 2y = 18 What is the x-value in the solution?

OpenStudy (anonymous):

3x + 4y = -10 (5x - 2y = 18)2 ----------------- 13x = 26 x = 2